RC meeting and Run Plan, Monday Oct. 23, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time: Monday - Friday at 3:30pm

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links


Hall A call-in instructions:

  • Call RC first.
  • Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
  • After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.

Hall Access procedure:

  • Access check list (binder on the counting room desk) needs to be filled out before making an access.
  • Scan (take pictures with phone) and post it to the logbook.
  • Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.

New instruction for shift crew when changing target:

  • Beam restoration instructions have been updated. Printed version is in the counting room.
  • When we change target, make sure the SOFs of Target BPMs are set to the correct values.
  • SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)

Status Update

Overall Summary


  • Weekend was very productive running, accumulated ~7.4 C of corrected charge since last RC meeting
  • Struggled with bleedthrough when Hall C went to 30 uA over the weekend - unresolved
  • Struggled with GEM trips because correlating with increased IC rates
    • MCC thinks this is due to orbit locks
    • Sounds like some work in planed for the maintenance day
    • MCC working on stop gap measures for the next 2 days
  • Sieve slit was inserted this morning around 8:40 for Optics data taking during beam studies
    • Running at 45 uA we saw ~2.6 kHz with a BBCal threshold of about 1.7 GeV
  • Plans for target laser work?
  • Getting encouragement from Doug to run at 50 uA

Shift Plan

  • Production Data


  • Up time has been very good this weekend

Opportunistically between spin ups (?):

  • H2 Reference data

Accelerator status

  • Maintenace Day Wednesday 10/25
  • If you want to get into the hall, please communicate with the RC to get this organized!

Data collected :

3 C of corrected charge collected between 7 am yesterday and today

Hall A Subsystems Status


  • Some failed fits on a 2 or 3 NMRs over the weekend
  • Polarization staying pretty stable around 46-47 % while in beam


  • HCal (Jiwan Poudel)
    • TDC Still looks good
  • Bigbite (Arun Tadepalli)
    • Mostly good (Provakar)
    • SC 18-19 missing
    • PreShower R-24 having issues, out of acceptance, will look at on maintenance day
  • DAQ (Alex Camsonne)
  • GEMs (Holly Szumila-Vance)
    • Mostly good (Andrew)
    • SBS Layer 2 hole
    • SBS LV hardware issue
  • Grinch
    • Rates too high?
  • Hodoscopes (Rachel Montgomery):
    • Looks fine (Gary)


  • SBS and BB are on.

Hall access requests

If anyone wants to make an access, let the RC know!

Shift schedule


  • None

Please check when your shifts are!


Please sign up for any open shifts!

Running Plan

  • Standard production running at 45uA, with NMRs every 3 hours.

==> List of items in case He3 production running isn’t ready:

  1. Take data on reference cell at 10 atm H2 [85% (4200 elastic events) more data required for Bigbite momentum calibration]
  2. BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on) (? is this different than the optics runs?)